OpenStudy (mrmanman305):

Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y = Two divided by x squared. + 3 please help someone

OpenStudy (jchick):

F(x) = 7/x g(x)=x²+10 Then f(g(x))= 7/(x²+10) By the way g(f(x))=(7/x)² + 10

OpenStudy (mrmanman305):

THANKS

OpenStudy (jchick):

No problem!